# Online Matroid Embeddings

### v1

@@ 6 Graphic and Regular Matroids
Assume that such embedding exists and let $G=(V,E)$ be the graph representing BigM. Now, consider two graphic matroids $\textbf{M}_{1}$ and $\textbf{M}_{2}$ represented respectively by the graphs in the left and right of Figure 2. Observe that when restricted to $\{a,b,c\}$, the matroids are identical since every non-empty subset of elements is independent. When an OMM observes the restriction to those three elements, it has not way to know in which matroid we are in, so it needs to map those three edges to the same edges of the graph $G$ representing BigM. Let’s denote those edges by $a^{\prime},b^{\prime},c^{\prime}$. Let also $d^{\prime},e^{\prime},f^{\prime}$ be the edges in $G$ that the edges of the left matroid are mapped to and let $g^{\prime}$ be the edge that the $g$ edge in the right matroid is mapped to. Note that $\{a^{\prime},b^{\prime},d^{\prime}\}$, $\{b^{\prime},c^{\prime},e^{\prime}\}$ and $\{a^{\prime},c^{\prime},f^{\prime}\}$ must form cycles in $G$. The only way that this is possible while keeping $\{a^{\prime},b^{\prime},c^{\prime}\}$ independent is if edges $a^{\prime},b^{\prime},c^{\prime}$ all share an endpoint. However, edges $a^{\prime},b^{\prime},c^{\prime},g^{\prime}$ must form a cycle in $G$ as well, which is not possible if the first three edges share an endpoint. ∎
$a$$b$$c$$d$$f$$e$$a$$b$$c$$g$
@@
If a host matroid BigM admits an online matroid embedding of all rank $n$ graphic matroids, then BigM must also admit an online matroid embedding of all rank $n$ binary matroids.
@@ 7 Laminar Matroids
[…]
@@
The OMM is defined inductively. If M is an empty matroid then $f_{\textbf{M},\pi}$ is the trivial morphism from the empty matroid to $\textbf{BigM}=\mathbb{F}^{n}$. Otherwise let $n=|\textbf{M}|$, let $(\textbf{M}^{\prime},\pi^{\prime})$ denote the restriction of $(\textbf{M},\pi)$ to $[n-1]$, let $g$ denote the morphism $f_{\textbf{M}^{\prime},\pi^{\prime}}$, and let $u=\pi(n)$. The morphism $f=f_{\textbf{M},\pi}$ is defined as follows. For $u^{\prime}\neq u$ we set $f(u^{\prime})=g(u^{\prime})$. If $u$ doesn’t belong to any circuit with previously arrived elements, define $f(u)$ to be any element outside the linear span of $f(\textbf{M}^{\prime})$. If $u$ is in some circuit, define $A\subseteq\textbf{M}$ to be the intersection of $\textsf{span}(C)$ over all circuits $C$ in M containing $u$. By Theorem 7.2, $A$ is the span of some circuit in M and hence a flat. Now, choose $f(u)$ to be an element in the linear subspace $V=\textsf{span}_{\mathbb{F}^{n}}(f(A\setminus\{u\}))$ that is not contained in $V_{S}=\textsf{span}_{\mathbb{F}^{n}}(f(S))$ for every subset $S$ of M such that $S\cup\{u\}$ is independent on M.
To justify that $f(u)$ is well-defined, we must argue that the linear subspace $V$ contains at least one element that is not in $V_{S}$ for every subset $S$ such that $S\cup\{u\}$ is independent. We will use a counting argument that consists of showing that $V_{S}\cap V$ is a proper linear subspace of $V$ for every such $S$, and then observing that a vector space over a field with at least $2^{n}$ elements cannot be expressed as the union of $2^{n}$ or fewer proper linear subspaces. To show that $V_{S}\cap V$ is a proper linear subspace of $V$ we argue by contradiction. Let $C$ be a circuit containing $u$ such that $A=\textsf{span}_{\textbf{M}}(C)$. As $S\cup\{u\}$ is independent, there must be some $u^{\prime}\in C\setminus\{u\}$ such that $S\cup\{u^{\prime}\}$ is also independent. Then, since $g:\textbf{M}^{\prime}\to\textbf{BigM}$ is a matroid morphism, $g(u^{\prime})$ is linearly independent of $g(S)$. In particular, $g(u^{\prime})\in V\setminus V_{S}$ cannot and hence $V_{S}\cap V$ is a proper linear subspace of $V$ as claimed.
@@
8 Matroids of Bounded Rank
@@
The matroid $\textbf{U}_{1,1}\oplus T$ has ground set $\{0,1\}$ and the only independent set is $\{1\}$. Give a matroid $\textbf{M}\in\mathcal{M}_{n,1}$ and an element $e$ of the ground set of M we map it to $1$ if $\textsf{rank}_{\textbf{M}}(\{e\})=1$ and to 00 otherwise. The mapping is constructed online since it only depends on the rank of a single-set element consisting of the element we are processing. It is also simple to check that it is a morphism since M has rank $1$. ∎
@@
The matroid $\textbf{U}_{n,2}\oplus T$ has ground set $\{0,1,2,\ldots,n\}$ and the independent sets are subsets of at most $2$ elements that don’t contain zero. To construct an embedding $f:\textbf{M}\rightarrow\textbf{U}_{n,2}\oplus T$, when we process an element $e$, we first check if $\textsf{rank}_{\textbf{M}}(\{e\})=0$. If so, we map it to 00. Otherwise, for every element $a$ processed before $e$, we check if $\textsf{rank}_{\textbf{M}}(\{a,e\})=1$. If so, we map $f(e)=f(a)$. Otherwise, we map $e$ to the first unused index in $[n]$. ∎
@@
Now suppose BigM is a matroid and we’re trying to construct an online embedding of a matroid M into BigM, where M is a rank-3 matroid that could either be $\textbf{M}_{1}$ or $\textbf{M}_{2}$. When the first six elements of M arrive, we have no way of distinguishing whether the input sequence is going to be $\textbf{M}_{1}$ or $\textbf{M}_{2}$. The online embedding algorithm chooses some function $h$ mapping $\{a,b,c,d,e,f\}$ to a six-element subset of BigM. Denote the images of $a,b,c,d,e,f$ in BigM by capital letters, for example $h(a)=A$. Now suppose there are two different extensions of $h$ to the domain ${a,b,c,d,e,f,g}$, denoted by $h_{1}$ and $h_{2}$, such that $h_{i}$ is an embedding of $\textbf{M}_{i}$ into BigM for each $i$. Let $G_{1}=h_{1}(g)$ and $G_{2}=h_{2}(G)$. Then the following must hold:
@@
which evalutes to $2+1-1=2$. This contradicts the fact that $\{E,F,G2\}$ has rank $3$. ∎
@@
$\begin{bmatrix}0&2&1&1&2&0&1\\ 1&1&2&0&0&2&1\\ 2&0&0&2&1&1&1\end{bmatrix}\quad\text{ and }\quad\begin{bmatrix}0&2&1&1&2&0&1\\ 1&1&2&0&0&1&1\\ 2&0&0&2&1&1&1\end{bmatrix}$
@@ 9 Approximate Embeddings
Theorem 9.2 (Abdolazimi et al[1], Dughmi et al [12]).
@@
As we discussed, constructing an $n$-approximate embedding is trivial. For binary matroids, we complement the lower bound of Dughmi et al [12] with an algorithm to construct a $O(n/\log n)$ embedding into the free matroid. Moreover, this embedding can be computed online.
@@
Open Questions.
We end with two open questions on approximate embeddings. The first is whether we can generalize the approximate embedding of Korula-Pál to the online setting. A possible approach is through the construction in Lemma 9.4. The embedding of Korula-Pál is a randomization between two embeddings. The first is exactly the embedding in Lemma 9.4 when the basis $b$ of the graphic matroid is a simple path in the underlying graph. The second can be obtained from a construction analogous to Lemma 9.4 taking the largest index instead of the smallest index. The challenge converting it to the online setting is identifying a ”good basis” in an online fashion knowing only the matroid structure but not the graph.
Open Question 1.

### v2

@@ 6.1 Graphic and Regular Matroids
Assume that such embedding exists and let $G=(V,E)$ be the graph representing BigM. Now, consider two graphic matroids $\textbf{M}_{1}$ and $\textbf{M}_{2}$ represented respectively by the graphs in the left and right of Figure 3. Observe that when restricted to $\{a,b,c\}$, the matroids are identical since every non-empty subset of elements is independent. When an OMM observes the restriction to those three elements, it has no way to know in which matroid we are in, so it needs to map those three edges to the same edges of the graph $G$ representing BigM. Let’s denote those edges by $a^{\prime},b^{\prime},c^{\prime}$. Let also $d^{\prime},e^{\prime},f^{\prime}$ be the edges in $G$ that the edges of the left matroid are mapped to and let $g^{\prime}$ be the edge that the $g$ edge in the right matroid is mapped to. Note that $\{a^{\prime},b^{\prime},d^{\prime}\}$, $\{b^{\prime},c^{\prime},e^{\prime}\}$ and $\{a^{\prime},c^{\prime},f^{\prime}\}$ must form cycles in $G$. The only way that this is possible while keeping $\{a^{\prime},b^{\prime},c^{\prime}\}$ independent is if edges $a^{\prime},b^{\prime},c^{\prime}$ all share an endpoint. However, edges $a^{\prime},b^{\prime},c^{\prime},g^{\prime}$ must form a cycle in $G$ as well, which is not possible if the first three edges share an endpoint. ∎
aabbccddffeeaabbccgg
@@
If a host matroid BigM is regular and admits an online matroid embedding of all rank $n$ graphic matroids, then BigM must also admit an online matroid embedding of all rank $n$ binary matroids.
@@ 6.2 Laminar Matroids
Denote the ground sets of $\textbf{M},\textbf{N}$ by $M,N$, respectively. If $\mathcal{A}$ is a laminar family of subsets of $N$ such that N is a laminar matroid with respect to the function $c:\mathcal{A}\to\mathbb{Z}_{+}$, then the family of sets $f^{-1}(\mathcal{A})$ consisting of the sets $f^{-1}(A)$ for each $A\in\mathcal{A}$ is a laminar family, and M is a laminar matroid with respect to the function $\tilde{c}:f^{-1}(\mathcal{A})\to\mathbb{Z}_{+}$ defined by
@@
The OMM is defined inductively. If M is an empty matroid then $f_{\textbf{M},\pi}$ is the trivial morphism from the empty matroid to $\textbf{BigM}=\mathbb{F}^{n}$. Otherwise let $n=|\textbf{M}|$, let $(\textbf{M}^{\prime},\pi^{\prime})$ denote the restriction of $(\textbf{M},\pi)$ to [n1][n-1], let $g$ denote the morphism $f_{\textbf{M}^{\prime},\pi^{\prime}}$, and let $u=\pi(n)$. The morphism $f=f_{\textbf{M},\pi}$ is defined as follows. For $u^{\prime}\neq u$ we set $f(u^{\prime})=g(u^{\prime})$. If $u$ doesn’t belong to any circuit with previously arrived elements, define $f(u)$ to be any element outside the linear span of $f(\textbf{M}^{\prime})$. If $u$ is in some circuit, define $A\subseteq\textbf{M}$ to be the intersection of $\textsf{span}(C)$ over all circuits $C$ in M containing $u$. By Theorem 6.5, $A$ is the span of some circuit in M and hence a flat. Now, choose $f(u)$ to be an element in the linear subspace $V=\textsf{span}_{\mathbb{F}^{n}}(f(A\setminus\{u\}))$ that is not contained in $V_{S}=\textsf{span}_{\mathbb{F}^{n}}(f(S))$ for every subset $S$ of M such that $S\cup\{u\}$ is independent on M.
To justify that $f(u)$ is well-defined, we must argue that the linear subspace $V$ contains at least one element that is not in $V_{S}$ for every subset $S$ such that $S\cup\{u\}$ is independent. We will use a counting argument that consists of showing that $V_{S}\cap V$ is a proper linear subspace of $V$ for every such $S$, and then observing that a vector space over a field with at least $2^{n}$ elements cannot be expressed as the union of $2^{n}$ or fewer proper linear subspaces. To show that $V_{S}\cap V$ is a proper linear subspace of $V$ we argue by contradiction. Let $C$ be a circuit containing $u$ such that $A=\textsf{span}_{\textbf{M}}(C)$. As $S\cup\{u\}$ is independent, there must be some $u^{\prime}\in C\setminus\{u\}$ such that $S\cup\{u^{\prime}\}$ is also independent. Then, since $g:\textbf{M}^{\prime}\to\textbf{BigM}$ is a matroid morphism, $g(u^{\prime})$ is linearly independent of $g(S)$. In particular, $g(u^{\prime})\in V\setminus V_{S}$ and hence $V_{S}\cap V$ is a proper linear subspace of $V$ as claimed.
@@
6.3 Matroids of Bounded Rank
@@
The matroid $\textbf{U}_{1,1}\oplus T$ has ground set $\{0,1\}$ and the only independent set is $\{1\}$. Given a matroid $\textbf{M}\in\mathcal{M}_{n,1}$ and an element $e$ of the ground set of M we map it to $1$ if $\textsf{rank}_{\textbf{M}}(\{e\})=1$ and to 0 otherwise. The mapping is constructed online since it only depends on the rank of a single-set element consisting of the element we are processing. It is also simple to check that it is a morphism since M has rank $1$. ∎
@@
The matroid $\textbf{U}_{n,2}\oplus T$ has ground set $\{0,1,2,\ldots,n\}$ and the independent sets are subsets of at most $2$ elements that don’t contain zero. To construct an embedding $f:\textbf{M}\rightarrow\textbf{U}_{n,2}\oplus T$, when we process an element $e$, we first check if $\textsf{rank}_{\textbf{M}}(\{e\})=0$. If so, we map it to 0. Otherwise, for every element $a$ processed before $e$, we check if $\textsf{rank}_{\textbf{M}}(\{a,e\})=1$. If so, we map $f(e)=f(a)$. Otherwise, we map $e$ to the first unused index in $[n]$. ∎
@@
Now suppose BigM is a matroid and we’re trying to construct an online embedding of a matroid M into BigM, where M is a rank-3 matroid that could either be $\textbf{M}_{1}$ or $\textbf{M}_{2}$. When the first six elements of M arrive, we have no way of distinguishing whether the input sequence is going to be $\textbf{M}_{1}$ or $\textbf{M}_{2}$. The online embedding algorithm chooses some function $h$ mapping $\{a,b,c,d,e,f\}$ to a six-element subset of BigM. Denote the images of $a,b,c,d,e,f$ in BigM by capital letters, for example $h(a)=A$. Now suppose there are two different extensions of $h$ to the domain a,b,c,d,e,f,g{a,b,c,d,e,f,g}, denoted by $h_{1}$ and $h_{2}$, such that $h_{i}$ is an embedding of $\textbf{M}_{i}$ into BigM for each $i$. Let $G_{1}=h_{1}(g)$ and $G_{2}=h_{2}(G)$. Then the following must hold:
@@
which evalutes to 2+11=22+1-1=2. This contradicts the fact that $\{E,F,G2\}$ has rank $3$. ∎
@@
[…]
@@ 7 Approximate Embeddings
Theorem 7.2 (Abdolazimi et al. [3], Dughmi et al. [17]).
@@
As we discussed, constructing an $n$-approximate embedding is trivial. For binary matroids, we complement the lower bound of Dughmi et al. [17] with an algorithm to construct a $O(n/\log n)$ embedding into the free matroid. Moreover, this embedding can be computed online.
@@
No Constant-Approximate Embedding from Graphical Matroid to Free Matroid
Next, we show that there is no constant approximate online embedding of a graphical matroid into a free matroid when the elements are revealed online with access to an independence oracle. This is in contrast to the result of [30] that allows us to construct an approximate offline embedding.
Theorem 7.7.